The sum of two consecutive even integers is at least nine more that half the sum of the next consecutive even integer and twice the first. Find the value of the first integer.
Answer by stanbon(75887)

The sum of two consecutive even integers is at least nine more that half the sum of the next consecutive even integer and twice the first. Find the value of the first integer.
-------
1st:: 2x
2nd: 2x+2
------
Equation:
4x+2 >= (1/2)[(2x+4)+ 2(2x)] + 9
----
4x+2 >= (1/2)[6x+4]+9
4x+2 >= 3x+11
---
x >= 9
------------------
Ans: 1st integer 2x >= 18
-----------------------
Cheers,
Stan H.
--------------